OpenStudy (anonymous):

Hi simple radical question but I just dont know how to solve √(360)-√(160)-√(90) ???

OpenStudy (anonymous):

that equals 6sqrt(10) - 4sqrt(10) - 3sqrt(10) = -sqrt(10)

OpenStudy (anonymous):

I don't want to be rude but I am still confused.

OpenStudy (anonymous):

\[\sqrt{360}=\sqrt{36*10}=\sqrt{36}*\sqrt{10}=6\sqrt{10}\] apply the same logic with the other two terms

OpenStudy (anonymous):

oh ok thank you so much lol Its clear now

OpenStudy (anonymous):

its all good!:)